OpenStudy (anonymous):

This is a question on my math homework, but I can't make any sense of it. Here it is: find the equation of the normal to the curve with equation y=(x^3)+1. Is there a typo or am I just missing something?

OpenStudy (anonymous):

The normal is a line perpendicular to the tangent line

OpenStudy (anonymous):

Could you explain how to do this problem? It has been a while since I have done this kind of problem.

OpenStudy (anonymous):

for what point are you trying to find the normal to?

OpenStudy (anonymous):

Also if the way I have put in the equation is confusing it should be: \[y=x^3+1\]

myininaya (myininaya):

thats what i was fixing to ask lol

OpenStudy (anonymous):

(1,2)

OpenStudy (anonymous):

k... now take the derivative of you function.... plug in the values (1,2) this will give you the slope, Let me know what you get

myininaya (myininaya):

find f'(x) and evaluate it at x=1 then take the negative reciprocal of that this gives you the slope of the normal line

OpenStudy (anonymous):

for the derivative I got \[y=3x^2\]

OpenStudy (anonymous):

and now I put in 2 for y and 1 for x, such that \[2=3(1)^2\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

no..... sorry just 1 for x, bc y' is the slope

OpenStudy (anonymous):

so then y=3

OpenStudy (anonymous):

and to get x i plug in 3 for y in the derivative?

OpenStudy (anonymous):

yeah thats your slop for the tanget, now you have to find the slope for the normal

OpenStudy (anonymous):

since the normal is perpendicular to the tangent, take the reciprocal of the slope of the tangent to get the slope of the normal

OpenStudy (anonymous):

So if the sope of the tangent is 3, then the slope of the normal is 1/3, correct?

OpenStudy (anonymous):

close... -1/3

OpenStudy (anonymous):

If the reciprocal is the number written as a fraction then flipped why did the sign change?

OpenStudy (anonymous):

sorry it is 1/3, I'm being retarded

OpenStudy (anonymous):

and then thats it? the answer is 1/3?

OpenStudy (anonymous):

now use the point slope formula to get the equation of the normal

OpenStudy (anonymous):

refresh my memory, what is the point slope formula? Or can i just use y=mx+b?

OpenStudy (anonymous):

wait the slope is -1/3 bc perpendicular slope have opposite signs and the point slope formula is \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

perpendicular lines are negative reciprocals

OpenStudy (anonymous):

ok. y1 and x1 are from the points given at the start, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and m=\[-1/3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so plugging everything in gets\[3-2=-1/3(1-1)\]

OpenStudy (anonymous):

wich solves to 1=0

OpenStudy (anonymous):

no.... you are looking y, the values you have are y1, x1, and m

OpenStudy (anonymous):

y1=2, x1=1 and m=-1/3

OpenStudy (anonymous):

so then it should be \[y-2=-1/3(x-1)\]

OpenStudy (anonymous):

yeah... now put it in terms of y=mx+b

OpenStudy (anonymous):

\[2=-1/3(1)+b\]

OpenStudy (anonymous):

and solve for b?

OpenStudy (anonymous):

no... so you have \[y-2=-\frac{1}{3}(x-1)\] Expand this and you get \[y-2=-\frac{1}{3}x+\frac{1}{3}\] now combine like terms \[y=-\frac{1}{3}x+\frac{1}{3}+2\] \[y=-\frac{1}{3}x+\frac{7}{3}\] which is in the form of y=mx+b

OpenStudy (anonymous):

could you also distribute the negative and move the 1/3 from \[y-2=-1/3(x-1)\] to give 3y-6=-x+1

OpenStudy (anonymous):

then move -x+1 over to give x+3y-7=0

OpenStudy (anonymous):

yes you could

OpenStudy (anonymous):

ok thank you for the help. yo just got a new fan!

OpenStudy (anonymous):

also now i'll use this site for help with math if the teacher is unavailable. Thanks again!

OpenStudy (anonymous):

Cheers , hope I didin't confuse you

OpenStudy (anonymous):

no you did fine. Have to go now bye